# Enhanced Single-Photon Detector: A framework for Superconducting-Level Performance without cryogenic cooling

### v2

@@ 2.4 Performance Simulations
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Table 3: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
@@
2.5 Application on QKD
@@
3 Discussion
We now discuss the technological feasibility, as well as limitations and avenues for future work.
3.1 Technological Feasibility and Performance under Degraded Parameters
3.1.1 State Preparations and Measurements
3.1.2 Controlled Gates
However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations. While the ESPD design remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded. Specifically, if the post-selection efficiency of the controlled gate is $N$, the average transmission rate from one effective module to the next will be $p^{\frac{1}{N}}$. Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate. Table 4 and Fig.4 demonstrate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.). This result indicates robustness, validating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
(a) DE $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
(b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
Figure 4: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
Table 4: Performance evolution of an ESPD initialized with a baseline detector from [28], with $p$ varied from 0.80 to 0.96. Experiments are run for $(n,k)=(5,2)$ on each level, and all other settings remain consistent with those described in Section 2.4.
@@ 3.1.2 Controlled Gates
We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38]. Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% [39, 40, 41, 42]. However, it is also worth testing the effectiveness of the design under degraded parameters to verify its suitability for standard fabrication processes with lower technology requirements. Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5 with visual representation in Fig.5, and $P=0.40$, (see Table 6 with visual representation in Fig.6). All other settings remain consistent with those described in Section 2.4. The numerical results confirm the robustness of the scheme: Comparison of Table 5 to Table 2 demonstrates that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This demonstrated robustness indicates that the technological demands of the ESPD method are fully compatible with existing non-superconducting quantum technologies.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
@@
3.1.3 Choice of initialized SPDs
It is also important to note that the requirements for SPD used for initialization are minimal, further lowering the barrier to implementation. As demonstrated in comparing Table 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, simulations confirm that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable initialization.
3.1.4 Auxiliary Sources
3.2 Limitations and Future Works
Despite the significance of the ESPD framework, some limitations remain, which are left for future investigations.
3.2.1 Theoretical Completeness
The theoretical framework established in Section 2.2 remains an area for deeper exploration. While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of such a non-linear and time-variant system and the sufficiency for practical technology employments by numerical designs, these specialized mathematical studies are reserved for the future.
3.2.2 Resource Overhead vs. Environmental Simplicity
While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels. Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections. Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD. For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (if not applying techniques such as multiplexing), resulting in hundreds of integrated components. While not posing a fundamental physical barrier (as discussed in Section 3.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
3.2.3 Optimization Design
The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments. Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$). Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence. As shown in Section 2.4, the optimal choice of parameters $n$ (number of auxiliary signals in each level) and $k$ (decision threshold) is crucial: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost. Furthermore, as shown in Table 7 and visualized in Fig.7, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
@@ 3.2.3 Optimization Design
Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
@@
The optimization problem also encompasses the optimal circuit topology. To mitigate cumulative error propagation, minimizing the maximal circuit length is essential. As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the design: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1. However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
3.2.4 Physical Integrations and Experiments
@@
Despite the convincing demonstration of effectiveness via numerical simulations in Section 2.4, physical implementations of the ESPD scheme require dedicated experimental validations. Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization. However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components. Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation. Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
3.2.5 Count Rate and Timing Jitter
@@
4 Conclusion
We have introduced the Enhanced Single-Photon Detector (ESPD) framework, a robust methodology for upgrading low-performance SPDs into devices whose capabilities can match or even surpass those of state-of-the-art superconducting SPDs. The distinguished advantage of the ESPD paradigm is its foundation on non-superconducting components entirely, obviating the demanding cryogenic operational and infrastructural requirements of SNSPDs, and thereby enabling high-performance SPDs to be utilized beyond specialized low-temperature laboratories. Our numerical analysis shows that a legacy non-superconducting SPD (with $\text{DE}\approx 59\%$ and $\text{DCR}\approx 10^{-2}$) can be enhanced to achieve $\text{DE}>93\%$ and $\text{DCR}<10^{-9}$, reducing the minimal tolerable channel transmission rate in quantum communication by several orders of magnitude. Although physically realizing such a large-scale architecture requires substantial integration and engineering efforts, the ESPD framework represents a paradigm shift, trading resource overhead for operation-environmental simplicity. It opens the door to high-performance quantum-optical applications without cryogenic cooling, providing a clear and viable path toward the widespread implementation of quantum optical tasks in condition-constrained environments.
@@ 4 Conclusion
Funding Statements
There is no funding for this work.
Author Contributions
Data and Code Availability

### v3

@@ II.4 Performance Simulations
Table 3: Performance of an ESPD initialized with a SPD from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
@@
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, we present numerical simulations for the ESPD paradigm, indicating that substantial performance gains are achievable even with moderate initial hardware specifications. For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $93\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
Simulation Parameters: We evaluated the performance of the ESPD scheme using two distinct baseline SPDs. The first represents an older technology, characterized by low performance parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section III.1 for parameter discussions). We also implemented additional simulations using degraded parameter settings to evaluate robustness (also see Section III.1).
Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig. 2 and 3. The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs. Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
II.5 Application on QKD
@@
Table 4: Performance of an ESPD initialized with a SPD from [28], with $p$ varied from 0.80 to 0.96. Experiments are run for $(n,k)=(5,2)$ on each level, and all other settings remain consistent with those described in Section II.4.
@@
III Discussion
In the absence of experimental implementation, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations and directions for future work.
III.1 Technological Feasibility and Performance under Degraded Parameters
III.1.1 State Preparations and Measurements
III.1.2 Controlled Gates
(a) DE $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
(b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
Figure 4: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section II.4.
@@ III.1.2 Controlled Gates
However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations. While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded. Specifically, if the post-selection efficiency of the controlled gate is $N$, the average transmission rate from one effective module to the next will be $p^{\frac{1}{N}}$. Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section II.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate. Table 4 and Fig.4 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.). This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
@@
We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38]. Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% [39, 40, 41, 42]. However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements. Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5 with visual representation in Fig.5, and $P=0.40$, (see Table 6 with visual representation in Fig.6). All other settings remain consistent with those described in Section II.4. The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This robustness indicates that the technological demands of the ESPD method are fully compatible with existing non-superconducting quantum technologies.
III.1.3 Choice of initialized SPDs
It is also important to note that the requirements for SPD used for initialization are minimal, further lowering the barrier to implementation. As supported in comparing Table 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, simulations suggest that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable initialization.
III.1.4 Auxiliary Sources
III.2 Limitations and Future Works
Despite the significance of the ESPD framework, some limitations remain and warrant future investigation.
III.2.1 Theoretical Completeness
The theoretical framework established in Section II.2 remains an area for deeper exploration. While the numerical designs presented in Section II.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of such a non-linear and time-variant system and the sufficiency for practical technology employments by numerical designs, these specialized mathematical studies are reserved for the future.
III.2.2 Resource Overhead vs. Environmental Simplicity
Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section II.4
@@ III.2.1 Theoretical Completeness
While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels. Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections. Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD. For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (if not applying techniques such as multiplexing), resulting in hundreds of integrated components. While not posing a fundamental physical barrier (as discussed in Section III.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
III.2.3 Optimization Design
The inherent exponential resource scaling discussed in Section III.2.2 necessitates optimization for future practical deployments. Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$). Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence. As shown in Section II.4, the optimal choice of parameters $n$ (number of auxiliary signals in each level) and $k$ (decision threshold) is crucial: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost. Furthermore, as shown in Table 7 and visualized in Fig.7, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
@@ III.2.3 Optimization Design
The optimization problem also encompasses the optimal circuit topology. To mitigate cumulative error propagation, minimizing the maximal circuit length is essential. As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section II.1. However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
III.2.4 Physical Integrations and Experiments
@@
Despite the convincing indication of effectiveness via numerical simulations in Section II.4, physical implementations of the ESPD scheme require dedicated experimental validations. Reassuringly, as established in Section III.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization. However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components. Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation. Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
III.2.5 Count Rate and Timing Jitter
@@
IV Conclusion
We have introduced the enhanced single-photon detector (ESPD) framework, which reformulates single-photon detection as an iteratively enhanced process based on integrating state preparations, controlled operations, projective measurements, and multi-copy analysis, rather than as a fixed device-level operation. Within this system-level theoretical framework, detector performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components. Numerical simulations indicate that, starting from a conventional non-superconducting SPD, the ESPD framework can substantially enhance DE while strongly suppressing DCR, yielding effective performance metrics comparable to those of state-of-the-art superconducting SPDs, and significantly relaxing the minimal tolerable channel transmission rate in quantum communication. The ESPD framework makes explicit the trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure with circuit integration complexity. While the experimental realization of large-scale ESPD architectures will require substantial engineering effort, the parameter studies presented in this work, grounded in the performance of existing photonic components, support the in-principle viability of the framework. In essence, this work establishes a system-level perspective on photon detection, showing that intrinsic device limitations can be transcended through iterative quantum processing and providing a principled foundation for the development of high-performance, room-temperature SPDs.
Data and Code Availability
Author Contributions